OpenStudy (anonymous):

Find the 5th term of the sequence, an=4(3)n-1. For a medal and fan

OpenStudy (anonymous):

Is it multiple choice or?

OpenStudy (anonymous):

4(3)*5-1=4(3)*4 4x(3*4) 4x81 =324 maybe?

OpenStudy (campbell_st):

just substitute n = 5 and calculate the value

OpenStudy (anonymous):

240 324 972 424